Ag-Coop is a large farm cooperative with a number of agriculture. related manufacturing and service divisions. As a cooperatiProduction is limited to the 750,000 kilowatt-hours monthly capacity of the dehydrator. Due to different chemical makeup, eacSelling expenses are 20 percent of the sales price The Retail Sales Division manager has complained that the prices charged bRequired A Required B Required C Assume that joint production costs including fixed overhe cost per pound of each product, inRequired A Required B Required Assume that joint production costs including fixed ove if sold through the cooperatives RetaiRequired A Required B Required Assume that joint production costs including fixed overhead ar operating profit under both Sch
Answer #1

Part A

Allocated cost per pound

$4.75

Per pound

Output Mix

Kilowatt-hour per pound

Kilowatt-hour per 100 pounds input

Greenup

50

32

1600

Maintane

30

20

600

Winterizer

20

40

800

3000

Maximum processing = 750,000 kwh / 3,000 kwh/100 pounds = 25000 pounds of input

Fixed cost allocation (81250/25000)

$3.25 per pound

Feedstock cost

1.50

Joint costs

$4.75 per pound

Part B

Allocated cost

Greenup

$4.97

Per pound

Maintane

$4.26

Per pound

Winterizer

$4.93

Per pound

Total joint cost incurred in processing 30,000 pounds of input = 81250+(25000*1.50) = $118750

Quantities of each product produced:

Greenup (25000*50%)

12500

Maintane (25000*30%)

7500

Winterizer (25000*20%)

5000

25000

(1) Sales Price/pound

(2) Selling Cost/pound (20% of Sales Price) (1)*20%)

(3) NRV/pound (1) – (2)

(4) Number of pounds

(5) Total NRV (3)*(4)

Greenup

10.50

2.10

8.40

12500

105000

Maintane

9.00

1.80

7.20

7500

54000

Winterizer

10.40

2.08

8.32

5000

41600

25000

$200600

Allocated cost

Greenup (118750*(105000/200600)/12500)

$4.97

Per pound

Maintane (118750*(54000/200600)/7500)

$4.26

Per pound

Winterizer (118750*(41600/200600)/5000)

$4.93

Per pound

Part C

Schedule A

Schedule B

Operating profit

$81850.00

$71368.04

profit under current production schedule A

Total net realizable value

200600

Less joint costs incurred

118750

$81850

Outputs under production schedule B

Output Mix

Kilowatt-hour per pound

Kilowatt-hour per 100 pounds input

Greenup

60

32

1920

Maintane

10

20

200

Winterizer

30

40

1200

3320

Maximum processing = 750,000 kwh / 3,320 kwh/100 pounds = 22590 pounds of input

Quantities of each product produced:

Greenup (22590*60%)

13554

Maintane (22590*10%)

2259

Winterizer (22590*30%)

6777

22590

Profit under production schedule B = (8.40*13554)+(7.20*2259)+(8.32*6777)-(1.50*22590) -81250 = $71368.04